consider an experiment in which 50.0 ml of 0.090 M Strontium nitrate was reacted with 45.0 ml of 0.100 M potassium iodate. calculate the number of moles of the reagent

13. (20 points) Given: Sr(NO): + 2 Kio, → Sr(10h + 2 KNO) NE214 38 1. う , Consider an experiment in which 50.0 mL of 0.090 M Strontium nitrate was reacted with 45.0 mL of0.100 moles of each reagent. SHOW ALL woRK(24 12 um iodate. Calculate the number of (1网 nglanO.onon. sr(N 03)2 t a,k103 → xs, (10) z f2 kN03. 3:53 56 0mL 45 Dmレ b.308 0. Which is limiting reagent? a Kl0s:,ル071F- 2. 3. Calculate theoretical yield of Strontium iodate? SHOW ALL WORK 1007 LL 32.S6 If your partner recovered 0.238 g of strontium iodate, what would be the percent yield? SHOW ALL WORK 33.543
Answer #1

1. Outline to solve the problem: molarity (M) of the solution = no. of moles of compound (mol) / volume (L)

i.e. The no. of moles of a compound = Its volume (L) * its molarity (mol/L)

According to the given data:

The no. of moles of Sr(NO3)2 = 50*10-3 L * 0.09 mol/L = 4.5*10-3 mol

The no. of moles of KIO3 = 45*10-3 L * 0.1 mmol/mL = 4.5*10-3 mol